Tracey drives to work at 50 mph and is 1 minute late. If she leaves at the same time and drives 55 mph, she'll be 1 minute early. How far does Tracey live from work?
Answer by mananth(16949)

let distance to work be x miles
speed = 50
time = x/50
...
speed increase to 55 mph
time = x/55
..
x/50 - x/55 = 2/60
(55x-50x)/50*55=1/30
5x/50*55=1/30
5x=50*55/30
/5
x= (50*55)/(5*30)
x= 18.3 miles to work
...
m.ananth@hotmail.ca